OpenStudy (alexh107):

Solve (28/x) - (7/x^2) = 7

jhonyy9 (jhonyy9):

like a first step what will be the common denominator ?

OpenStudy (alexh107):

Um I'm not sure.

jhonyy9 (jhonyy9):

for x and x^2 what is the GCF ?

OpenStudy (alexh107):

x?

jhonyy9 (jhonyy9):

sorry not is this the right question for common denominator

jhonyy9 (jhonyy9):

the great common divisor ?

OpenStudy (alexh107):

Sorry I don't think I know

jhonyy9 (jhonyy9):

why not this is very easy for x^2 is x^2 yes ? bc. x^2 /x^2 = 1 and for x is too x^2 bc. x^2 /x = x ok. ?

OpenStudy (alexh107):

Okay

jhonyy9 (jhonyy9):

so and in this way you get that the common denominator is x^2

OpenStudy (alexh107):

Yes

jhonyy9 (jhonyy9):

so than x^2 is the common denominator with what will need multiplie the first fraction ?

OpenStudy (alexh107):

x?

jhonyy9 (jhonyy9):

yes sure

jhonyy9 (jhonyy9):

and on the other side the 7 ?

OpenStudy (alexh107):

So the first fraction multiplied would be 28x/x^2 and then the 7 would be 7x?

jhonyy9 (jhonyy9):

no for 7x so than the common denominator is x^2 and the 7 on the right side has the denominator 1 so than with what will be necessary multiplie the 7 ?

OpenStudy (alexh107):

x^2

jhonyy9 (jhonyy9):

yes sure exactly this is easy and you know it just more courage please ok. ?

jhonyy9 (jhonyy9):

so now please make what you write above

OpenStudy (alexh107):

28x/x^2 - 7/x^2 = 7/x^2 I'm confused about the 7/x^2 part though

jhonyy9 (jhonyy9):

no on the right side multiplie the 7 not just divide by x^2

OpenStudy (alexh107):

7x^2

jhonyy9 (jhonyy9):

yes so how will be ?

OpenStudy (alexh107):

28x/x^2 - 7/x^2 = 7x^2

OpenStudy (campbell_st):

I think the easy thing to do is multiply every term by x^2 and that will eliminate all denominators' then looking at things divide every term by 7 then I'd expect it's general quadratic formula for the solutions.

OpenStudy (campbell_st):

like this \[\frac{28}{x} \times x^2 - \frac{7}{x^2} \times x^2 = 7 \times x^2\]

OpenStudy (alexh107):

Ah, okay that makes sense.

OpenStudy (alexh107):

And then after that would I just use the quadratic formula to solve?